OpenStudy (anonymous):

write tan70-tan30/1-tan70tan30 as a function of a single angle, do not evaluate

OpenStudy (yrelhan4):

Hmm. are you sure this is the question? It would be easy if there's a plus sign in the denominator or the numerator? :P

OpenStudy (anonymous):

yes this is the question and i know trust me but unfortunately this is the question

OpenStudy (yrelhan4):

I would have to think. Not in the mind frame. Sorry.

OpenStudy (anonymous):

haha its okay

OpenStudy (anonymous):

you think i can use the sum and difference formula for this

OpenStudy (anonymous):

I have no idea......I was thinking just as yrelhan4.......

OpenStudy (anonymous):

ya but i was thinking maybe we use the tan(u+-v)

OpenStudy (anonymous):

yes.....that is true....let's try that and see what we would get.....

OpenStudy (anonymous):

i mean wouldn't i get tan(70+30)

OpenStudy (anonymous):

i mean i know it looks easy but i dont think its that easy idk

OpenStudy (anonymous):

i am on it.....

OpenStudy (anonymous):

\[\tan(u+v)=tanu+tanv/1-tanutanv\]

OpenStudy (anonymous):

yes....that expansion of tan is right...

OpenStudy (anonymous):

ya so i think thats the answer because the questions says do not evaluate

OpenStudy (anonymous):

if that should be the answer, the numerator should be +

OpenStudy (anonymous):

no it doesn't need to be cuz look at the expansion it has a negative in the numerator

OpenStudy (anonymous):

i am getting confused.....

OpenStudy (anonymous):

haha its okay

OpenStudy (anonymous):

hold on.....i am trying it...would post it

OpenStudy (anonymous):

\[\frac{ \tan70-\tan30 }{ 1-\tan70\tan30 }=\frac{ -(\tan-70+\tan30) }{ 1-\tan70\tan30 }\]

OpenStudy (anonymous):

i am not sure. it may contain mistakes.....i just want to give you a clue....